# Single-Pass Document Scanning for Question Answering

### v1

@@ A.3 Validation Sets Statistics
docfinQA $100$ 142 328142328142\,328142 328
muld_CAC $100$ 45 5204552045\,52045 520
ELITR_Bench $010$ 3281032810\,32810 328
narrativeqa $100$ 71 8437184371\,84371 843
qasper $100$ 5 27452745\,2745 274
@@
hotpotqa $001$ $31$\,3131 313
musique $100$ 2 23022302\,2302 230
@@ B.4 Synthetic Data Quality Evaluation
We provide and evaluate a few synthetic data examples generated from different synthetic data strategies in Tables 18, 19, and 20.
@@
The first example in 20 involves a question about an event that prompted an inquiry regarding a specific time during the group’s evening activities. Context 1 effectively answers this question by discussing these activities. However, Context 2, which frequently uses keywords like “evening” and “I,” creates a high semantic similarity with Context 1. Despite this similarity, Context 2 does not talk about the same event as Context 1 and does not contribute to answering the question in any sense. Similarly, the second example inquires about Thomas’s motivation for confessing. Context 1 clearly explains that Thomas confessed because he felt sorry for Mary. In contrast, Context 2 is unrelated to the question; it only contains negative words that might have some semantic similarity to the question.
Table 18: Linked-based Synthetic Data Examples
@@
Table 19: Chunk-based Synthetic Data Examples
@@ 6.1 EFFECTIVE DATE AND GRANT PERIOD
Table 20: Pair-based Synthetic Data Examples
@@ E.5 Ablation for Training Document Length
Table 25: Ablation study for the training document sequence length.
@@ E.7 Retrieval Comparison between Single-Pass Scanner and NV-Embed-v2
We demonstrate that the Single-Pass Scanner retrieves more relevant context than the embedding model by comparing the sentences retrieved by SPScanner 1.3B and NV-Embed-v2-7B for the following data points in the test set (examples where NV-Embed-v2-7B retrieves 50 sentences are in Tables 26, 27; examples where NV-Embed-v2-7B retrieves 5 chunks are in Tables 28).
In Table 26, NV-Embed-v2-7B successfully retrieves the semantically relevant sentence “Uh, like a test of availability,” which aligns with the query about “test components’ availability.” However, NV-Embed-v2-7B failed to retrieve a crucial follow-up sentence identifying PERSON 7 as the individual responsible for the test. This limitation highlights that, while NV-Embed-v2-7B effectively identifies phrases with high semantic similarity, it failed to capture broader contextual relationships needed for comprehensive information retrieval. In contrast, the SPScanner 1.3B demonstrated a stronger contextual understanding by successfully retrieving the sentence “So that’s another thing that, that [PERSON7], uh, uh, uh, should $<$unintelligible$>$ on,” which was crucial for fully answering the question.
@@
That we would know, uh, which of the parts of the pipeline are, performing badly in terms of translation quality.Uh, I just, uh-.It just occurred to me that there should be one more compilation target. And that would be like probing whether the components of the pipeline are up and running.Uh, like a test of availability. So that’s another thing that, that [PERSON7], uh, uh, uh, should $<$unintelligible$>$ on-. if you could put this on, uh, on the to do list or on the enhancement options, that would also be very useful.Uh, and another thing would be, uh, like live debugging, uh, of a of a pipeline such a speed of, uh, of that.Okay.And, uh, yes, and uh, so, and then the second item you have in your list [PERSON7].Please comment on that.(PERSON7) All, right.So, uh, next Friday, uh, like, next week somewhere there there is going to be a conference about [PROJECT13] and we are going to provide life subtitles and transcription.And because we will have some non native English speakers in there, so we will need to get some feedback, from the people that are using our subtitles.Preferab-, refe-, preferably life.So we can, uh, see a moments like, uh, where it was working and moments where it was not working.So, uh ,I will make some, uh, quick took-.(PERSON13) [PERSON15] already has such simple tool that you could adapt.Um, what is more-.What is missing is, uh, description, of, ah, like how to use the tool.And also more like a generic description of how people should look at the outputs.So, uh, it would be best, if you could get in touch with [PERSON18].Because I’ve asked [PERSON18] to like handle the soft, uh, soft things with the participants and also with the organisers.And, uh, um, you and [PERSON18] should prepare very simple instructions that the participants could follow. That we would know, uh, which of the parts of the pipeline are, performing badly in terms of translation quality.Uh, I just, uh-.It just occurred to me that there should be one more compilation target. And that would be like probing whether the components of the pipeline are up and running.Uh, like a test of availability. So that’s another thing that, that [PERSON7], uh, uh, uh, should $<$unintelligible$>$ on-.. if you could put this on, uh, on the to do list or on the enhancement options, that would also be very useful.Uh, and another thing would be, uh, like live debugging, uh, of a of a pipeline such a speed of, uh, of that.Okay.And, uh, yes, and uh, so, and then the second item you have in your list [PERSON7].Please comment on that.(PERSON7) All, right.So, uh, next Friday, uh, like, next week somewhere there there is going to be a conference about [PROJECT13] and we are going to provide life subtitles and transcription.And because we will have some non native English speakers in there, so we will need to get some feedback, from the people that are using our subtitles.Preferab-, refe-, preferably life.So we can, uh, see a moments like, uh, where it was working and moments where it was not working.So, uh ,I will make some, uh, quick took-.(PERSON13) [PERSON15] already has such simple tool that you could adapt.Um, what is more-.What is missing is, uh, description, of, ah, like how to use the tool.And also more like a generic description of how people should look at the outputs.So, uh, it would be best, if you could get in touch with [PERSON18].Because I’ve asked [PERSON18] to like handle the soft, uh, soft things with the participants and also with the organisers.And, uh, um, you and [PERSON18] should prepare very simple instructions that the participants could follow.

### v2

@@ A.3 Validation Sets Statistics
docfinQA $100$ 142 328142\,328
muld_CAC $100$ 45 52045\,520
ELITR_Bench $100$ 10 32810\,328
narrativeqa $100$ 71 84371\,843
qasper $100$ 5 2745\,274
@@
hotpotqa $100$ 1 3131\,313
musique $100$ 2 2302\,230
Table 8: Dataset statistics for validation set.
@@ B.4 Synthetic Data Quality Evaluation
We provide and evaluate a few synthetic data examples generated from different synthetic data strategies in Tables 19, 20, and 21.
@@
The first example in 21 involves a question about an event that prompted an inquiry regarding a specific time during the group’s evening activities. Context 1 effectively answers this question by discussing these activities. However, Context 2, which frequently uses keywords like “evening” and “I,” creates a high semantic similarity with Context 1. Despite this similarity, Context 2 does not talk about the same event as Context 1 and does not contribute to answering the question in any sense. Similarly, the second example inquires about Thomas’s motivation for confessing. Context 1 clearly explains that Thomas confessed because he felt sorry for Mary. In contrast, Context 2 is unrelated to the question; it only contains negative words that might have some semantic similarity to the question.
@@
Table 19: Linked-based Synthetic Data Examples
@@ 6.1 EFFECTIVE DATE AND GRANT PERIOD
Table 20: Chunk-based Synthetic Data Examples
@@
Table 21: Pair-based Synthetic Data Examples
@@ C.2 Multiple Choice Question Question-Answer Judging Prompt
C.3 Relevant Sentence Annotation Prompt
See Figure 12 for the relevant sentence annotation prompt.
Figure 12: Prompt for annotating relevant sentences for the given query.
@@ E.5 Ablation for Training Document Length
Table 26: Ablation study for the training document sequence length.
@@ E.7 Retrieval Comparison between Single-Pass Scanner and NV-Embed-v2
We demonstrate that the Single-Pass Scanner retrieves more relevant context than the embedding model by comparing the sentences retrieved by SPScanner 1.3B and NV-Embed-v2-7B for the following data points in the test set (examples where NV-Embed-v2-7B retrieves 50 sentences are in Tables 28, 29; examples where NV-Embed-v2-7B retrieves 5 chunks are in Tables 30).
In Table 28, NV-Embed-v2-7B successfully retrieves the semantically relevant sentence “Uh, like a test of availability,” which aligns with the query about “test components’ availability.” However, NV-Embed-v2-7B failed to retrieve a crucial follow-up sentence identifying PERSON 7 as the individual responsible for the test. This limitation highlights that, while NV-Embed-v2-7B effectively identifies phrases with high semantic similarity, it failed to capture broader contextual relationships needed for comprehensive information retrieval. In contrast, the SPScanner 1.3B demonstrated a stronger contextual understanding by successfully retrieving the sentence “So that’s another thing that, that [PERSON7], uh, uh, uh, should <<unintelligible>> on,” which was crucial for fully answering the question.
@@
However, overall outperforms SP-Scanner on conversational datasets. NV-embed’s superior performance on conversational datasets can be attributed to the specific nature of these tasks, which often do not require capturing long-range dependencies from distant parts of the document. Since embedding models can retrieve local context but falter at long-range dependencies, these conversational datasets avoid the weakness of embedding models. Of the 707 data points in the conversational datasets, 400 focus specifically on speaker identification tasks, divided into two categories: 200 questions about masked name entities and 200 requiring identification of speakers for given sentences. Both tasks only require knowing the immediate context to answer the question. We provide one example for each task in Table 27 below.
Example 1: Masked Name Entity
Task Masked Name Entities
Question Which character is \$\$MASK\$\$?
Context PATRICK: I do not care for your language. LEONARD: I’m only trying to help you guys. Leonard stops exercising. \$\$MASK\$\$: (CONT’D) You’re fat dad. Mom is fat. Us kids are fat. (CONTINUED). CONTINUED: PATRICK: What’s your point?
Answer LEONARD
Example 2: Speaker Identification
Task Speaker Identification
Question ‘Who would like to come in on that? Go on, Catherine.’ said by ___
Context “And Catherine, I just wonder, extending Alun’s point, …” said by Huw Irranca-Davies AM “On the point about being boring, generally boring is good as far as the EU is concerned …” said by Alun Davies AM “Who would like to come in on that? Go on, Catherine.” said by ___
Answer Huw Irranca-Davies AM
Table 27: Examples of retrieval tasks
@@
That we would know, uh, which of the parts of the pipeline are, performing badly in terms of translation quality.Uh, I just, uh-.It just occurred to me that there should be one more compilation target. And that would be like probing whether the components of the pipeline are up and running.Uh, like a test of availability. So that’s another thing that, that [PERSON7], uh, uh, uh, should <<unintelligible>> on-. if you could put this on, uh, on the to do list or on the enhancement options, that would also be very useful.Uh, and another thing would be, uh, like live debugging, uh, of a of a pipeline such a speed of, uh, of that.Okay.And, uh, yes, and uh, so, and then the second item you have in your list [PERSON7].Please comment on that.(PERSON7) All, right.So, uh, next Friday, uh, like, next week somewhere there there is going to be a conference about [PROJECT13] and we are going to provide life subtitles and transcription.And because we will have some non native English speakers in there, so we will need to get some feedback, from the people that are using our subtitles.Preferab-, refe-, preferably life.So we can, uh, see a moments like, uh, where it was working and moments where it was not working.So, uh ,I will make some, uh, quick took-.(PERSON13) [PERSON15] already has such simple tool that you could adapt.Um, what is more-.What is missing is, uh, description, of, ah, like how to use the tool.And also more like a generic description of how people should look at the outputs.So, uh, it would be best, if you could get in touch with [PERSON18].Because I’ve asked [PERSON18] to like handle the soft, uh, soft things with the participants and also with the organisers.And, uh, um, you and [PERSON18] should prepare very simple instructions that the participants could follow. That we would know, uh, which of the parts of the pipeline are, performing badly in terms of translation quality.Uh, I just, uh-.It just occurred to me that there should be one more compilation target. And that would be like probing whether the components of the pipeline are up and running.Uh, like a test of availability. So that’s another thing that, that [PERSON7], uh, uh, uh, should <<unintelligible>> on-.. if you could put this on, uh, on the to do list or on the enhancement options, that would also be very useful.Uh, and another thing would be, uh, like live debugging, uh, of a of a pipeline such a speed of, uh, of that.Okay.And, uh, yes, and uh, so, and then the second item you have in your list [PERSON7].Please comment on that.(PERSON7) All, right.So, uh, next Friday, uh, like, next week somewhere there there is going to be a conference about [PROJECT13] and we are going to provide life subtitles and transcription.And because we will have some non native English speakers in there, so we will need to get some feedback, from the people that are using our subtitles.Preferab-, refe-, preferably life.So we can, uh, see a moments like, uh, where it was working and moments where it was not working.So, uh ,I will make some, uh, quick took-.(PERSON13) [PERSON15] already has such simple tool that you could adapt.Um, what is more-.What is missing is, uh, description, of, ah, like how to use the tool.And also more like a generic description of how people should look at the outputs.So, uh, it would be best, if you could get in touch with [PERSON18].Because I’ve asked [PERSON18] to like handle the soft, uh, soft things with the participants and also with the organisers.And, uh, um, you and [PERSON18] should prepare very simple instructions that the participants could follow.
@@
E.8 Performance Change around 128-256k Tokens
As shown in Figure 4 of the main text, SPScanner’s performance unexpectedly improves at 128-256k tokens before declining linearly afterward. This counterintuitive pattern occurs because different context length intervals contain distinct document types. In Table 32 below, we show which document types appear at each context length range.
Token Range Description and Example Datasets
1–8k tokens Scientific research papers (qasper, paperqa); legal and nonfiction documents (multifieldqa_en); historical records (2wikimqa, altqa); meeting transcripts focused on speaker identification.
8–16k tokens Wikipedia passages and screenplays (hotpotqa); government documents (musique); lecture transcripts with multiple-choice questions (coursera); meeting transcripts with relationship queries; historical QA from Wikipedia (altqa_16k).
16–32k tokens Novels (narrativeqa); legal contracts (legal_contract_qa); passages combined from novels, historical texts, scientific forms, and dictionaries (loogle_CR_mixup_16k, loogle_MIR_mixup_16k, multifieldqa_en_mixup_16k).
32–64k tokens Combined passages from varied sources (loogle_CR_mixup_32k, loogle_MIR_mixup_32k, multifieldqa_en_mixup_32k); narrative tasks with villain–hero identification (muld_CAC).
64–128k tokens Combined passages from varied sources (loogle_CR_mixup_64k, loogle_MIR_mixup_64k, multifieldqa_en_mixup_64k); dialogue transcripts requiring character identification across long conversations (longdialogue_qa_eng).
128–256k tokens Novels with QA or multiple-choice formats (longbook_qa_eng, longbook_choice_eng); financial filings and reports (docfinQA); combined passages from varied sources (loogle_CR_mixup_128k, loogle_MIR_mixup_128k, multifieldqa_en_mixup_128k).
256k+ tokens Combined passages from novels, historical documents, scientific texts, and dictionaries (loogle_CR_mixup_256k, loogle_MIR_mixup_256k, multifieldqa_en_mixup_256k).
Table 32: Overview of Datasets by Context Length
E.9 SPScanner Qualitative Error Analysis
In this section, we provide a qualitative error analysis where GPT-4o Full-Context answers correctly but SPScanner 1.3B’s downstream generator answers incorrectly. In most cases, SPScanner fails because it only retrieves fragmented sentences. Table 33 presents three example questions along with their corresponding answers, SPScanner responses, and GPT-4o Full-Context responses.
Example 1: Dialogue QA (‘longdialogue_qa_eng‘)
Question Which character is \$\$MASK\$\$?
Answer CHRISTY
SPScanner Answer Lisa
GPT-4o Full-Context Answer Christy
Example 2: Narrative QA (‘longbook_qa_eng‘)
Question Why does Sir Dilan loose two boots in London?
Answer Angelo needed the scent from an old boot to lure his hound
SPScanner Answer Sir Dilan loses two boots in London because they were stolen and used in a plot against him, involving setting a hound on his track.
GPT-4o Full-Context Answer Sir Dilan loses two boots in London because Angelo needed an article of his attire to set the hound on his track, leading to the theft of his boots.
Example 3: Document Analysis (‘loogle_CR_mixup_16k‘)
Question Please select architectures are affected by the Baroque style according to the passage?
1. The Church of Saints Cosme and Damiano.
2. The Church and Convent of São Francisco in Salvador.
3. Portuguese military buildings before 16th century.
Answer 1, 2
SPScanner Answer 1
GPT-4o Full-Context Answer 1, 2
Table 33: Qualitative Analysis of SPScanner vs. Full-Context GPT-4o.
For the first example, SPScanner correctly retrieved the target sentence “\$\$MASK\$\$: (coyly) What a good friend” but missed the crucial preceding context: “CHRISTY: Don’t worry.He’s in good hands.Trent cracks the door and peers through.The light is dim, but he can make out that they’re starting to neck.He closes the door, satisfied.” Without this speaker tag establishing Christy’s presence in the sleeping compartment, the downstream generator selected the nearest visible female name (Lisa) from the retrieved sentences and answered incorrectly. In contrast, GPT-4o had access to the full contiguous passage showing the complete sleeping-compartment context where only Christy and Trent were present, enabling correct identification.
For the second example, SPScanner’s answer was less precise because while its retrieved sentences provided clues about stolen boots being used to set a hound on someone’s track, the final answer failed to explicitly identify the antagonist or explain that an old boot was needed specifically for its scent. GPT-4o directly identifies the complete sequence: stealing a new boot, finding it useless for scent-tracking, then stealing an old one that would carry the target’s scent. This led to a more complete explanation of why two boots were lost rather than just one.
For the third example, SPScanner failed because it only retrieved fragmented sentences that mentioned baroque features but lacked the complete contextual descriptions needed to properly identify which specific buildings were affected by the Baroque style. While the retrieved sentences contained references like “although the tower is partly baroque” and mentions of baroque ornamentation, the generator couldn’t definitively connect these features to both churches.